OpenStudy (anonymous):

At Mark's birthday party, Michael ate 3 pieces of pizza, tony ate 2 pieces, Stacey ate 4 pieces, and Larry ate 3 pieces. Now one fourth of the pizza is left. Into how many pieces was the pizza originally cut?

OpenStudy (anonymous):

i need help!

OpenStudy (anonymous):

3+2+4+3=(3/4)x Solving for x will give you the answer

OpenStudy (anonymous):

is it 13?

OpenStudy (anonymous):

No... 12=3/4x 16=x

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

Yes, 16 pieces originally